OpenStudy (anonymous):

Simplify 10x/5x+3 + 6/10x+6

OpenStudy (anonymous):

look: 10x+6=2(5x+3) so, 10x/(5x+3) +6/(2*(5x+3)) simplify second expression (divide by 2): 10x/(5x+3) +3/(5x+3) = (10x+3)/(5x+3)